OpenStudy (anonymous):

The length of one side of the regular pentagon is 18.3cm. What is the perimeter of the pentagon? GIVE MEDALS!

OpenStudy (ipwnbunnies):

How many sides does a pentagon have?

OpenStudy (ipwnbunnies):

It has 5 sides. A regular pentagon will have 5 sides of the same length. The perimeter of a shape is the length of the lines that make the shape. So, the perimeter is the sum of the lengths of all the sides. Since a pentagon has 5 sides, its perimeter will be 5*18.3cm = 18.3+18.3+18.3+18.3+18.3 = ?

OpenStudy (ipwnbunnies):

@Marzia123